OpenStudy (nathanjhw):

Suppose f(x)= sin (pi cos x) .On any interval where the inverse function y = f –1(x) exists, the derivative of f –1(x) with respect to x is:

OpenStudy (perl):

$$ \Large f(x) = \sin (\pi \cos(x)) $$

OpenStudy (perl):

is that the correct f(x)

OpenStudy (nathanjhw):

yes

OpenStudy (perl):

$$ \Large { y = \sin (\pi \cos( x) ) \\ \therefore \\ x = \sin (\pi \cos (y) ) \\ \implies \\ \arcsin(x) = \pi \cos( y) \\ \cos y = \frac{1}{\pi} \arcsin(x) \\ \implies\\ \\ y = \arccos (\frac{1}{\pi} \arcsin(x)) \\ \implies\\ \frac{dy}{dx}=- \frac{1}{\sqrt{1-(\frac{1}{\pi} \arcsin(x))^2~})}\cdot ... } $$

OpenStudy (perl):

is this a multiple choice question

OpenStudy (nathanjhw):

The answers are: -1/cos(pi cosx) where x and y are related by the equation (satisfy the equation) x= sin(pi cosy) -1/ pi sin x cos (pi cos x) where x and y are related by the equation x= sin(pi cosy) -1/ pi sin y cos (pi cos y) where x and y are related by the equation x= sin(pi cosy) -1/ cos (pi cos y) where x and y are related by the equation x= sin(pi cosy) -1/sin y cos (pi cos y) where x and y are related by the equation x= sin(pi cosy)

OpenStudy (nathanjhw):

@perl

OpenStudy (perl):

there is a trick we can use

OpenStudy (perl):

$$ f(x) = \sin (\pi \cos(x))\\ \LARGE \frac{dx}{dy} = \frac{1}{\frac{dy}{dx}} \\ \Large { \\ \therefore \\ \frac{dx}{dy} = \frac{1}{\cos (\pi \cos(x))\cdot \pi \cdot (-\sin(x)} \\ } $$

OpenStudy (nathanjhw):

So then it would be the second one.

OpenStudy (perl):

we have to be careful, this is the inverse with respect to y

OpenStudy (nathanjhw):

So then?

OpenStudy (perl):

i see your choices, some are in terms of y, some in terms of x

OpenStudy (nathanjhw):

Is it x or y?

OpenStudy (perl):

I believe y

OpenStudy (perl):

you can test this supposition with y = x^3 you know the inverse is y = x^(1/3) and the derivatives of these

OpenStudy (perl):

$$ \Large Given ~~ f(x) = \sin (\pi \cos(x))\\ \LARGE \frac{dx}{dy} = \frac{1}{\frac{dy}{dx}} \\ \Large { \\ \therefore \\ \frac{dx}{dy} = \frac{1}{\cos (\pi \cos(x(y)))\cdot \pi \cdot (-\sin(x(y)))} \\ \\ \therefore \\ ~\\ (f^{-1}(x) )' = \frac{1}{\cos (\pi \cos(y(x)))\cdot \pi \cdot (-\sin(y(x)))} \\ (f^{-1}(x) )' = \frac{1}{\cos (\pi \cos(y))\cdot \pi \cdot (-\sin(y))} \\ } $$

OpenStudy (nathanjhw):

Alright thanks!